OpenStudy (anonymous):

a man loses 24% of his money and after spending 75 % of the remainder, he left with Rs133, how much he had first ?

OpenStudy (anonymous):

Let him have y initially If he loses 24% hes left with 76% so let x=76y/100 (money left first) He spends 75% of x, so he has only 25% left so 25x/100 = 133

OpenStudy (anonymous):

so x = 532 y = 100x/76 100 x 532/76

OpenStudy (anonymous):

y=700...So he had 700 $ or Rs or whatever

OpenStudy (anonymous):

k, thnxx !

OpenStudy (anonymous):

you can also think in reverse way. Rs 133 is equal to 25% of what he had after losing. so 133*4 = 532 is the money he totally had after losing. And Rs 532 is 76%. So initially he had, \[\left( 532*100 \right)\div76\] = 700